Pure acid is to be added to a 10% acid solution to obtain 90L of 27% solution. How many Liters of 100% pure acid should be used to make the solution
Answer by lwsshak3(11628)

Pure acid is to be added to a 10% acid solution to obtain 90L of 27% solution. How many Liters of 100% pure acid should be used to make the solution.
***
let x=amt of pure acid to be used
90-x=amt of 10% acid solution
..
100%x+10%(90-x)=27%*90
x+9-.1x=24.3
.9x=15.3
x=17
amt of pure acid to be used=17 Liters